It takes mr Hack 15 minutes to bake some cupcakes and it takes mr babbit 20 minutes to bake the same amount of cupcakes. The cupcakes all have to be yellow cake with white icing and then rainbow sprinkles on top of the icing. How long will it take them if they both work together to make some cupcakes?
Answer by checkley77(12844)

15*20/(15+20)=300/35=8.57 min. working together to bake a bath of cupcakes.